kelly telephoned brook about a homework problem. Kelly said Four plus three times two is 14,isnt it? brook replied, No its 10. did someone make a mistake?can you explain where these two answer came from?
Answer by checkley77(12569)

4+3*2=4+6=10 no parenthesis.
(4+3)2=7*2=14 with parenthesis.